Prev Next

SW&IT Meeting Minutes 2020-05-04

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200420
LR20200420 - Bec, Matthieu 80.0
LR20200420 - Holmberg, Carl 80.0
LR20200420 - Edmeades, David 40.0
LR20200420 - Sola, Igor 79.5
LR20200420 - Hooper, Stephen 80.0
LR20200420 - Zhang, Yang 79.0
LR20200420 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • tcs test status and next action
  • fedora 31
  • AO training sessions - 3:00-4:30 daily this week
Dear All,

Starting today, and extending through Friday, May 8th, Doug, Jenny, Shane, Greg, and I will be conducting a five day course of instruction covering the LBT adaptive optics (AO) system.  The course is organized as follows:
Day 1 - An introduction to adaptive optics (Al)
Day 2 - FLAO - The LBT AO system for LUCI (Doug)
Day 3 - FLAO continued [Doug] + AO operational web pages (Greg)
Day 4 - LBTI AO - Differences from FLAO (Jenny)
Day 5 - The AO On-Sky Check-Out Program (Shane)
Each session will take place from 3:00 to 4:30 on zoom5:
zoom.us/j/2668361325
Course material can be found on the LBTO wiki here:
https://wiki.lbto.org/AdaptiveOptics/RoundTwoTraining
All LBTO staff members are invited to attend.

We look forward to seeing you today at 3:00 PM.

Best regards,

Al

Team reports

MB

  • Last week
    • OC, Management
    • phone interview, new ad for senior
    • git renaming for AO-supervisor/confcalib
    • DMS wip satellite imagem, selenium script to extract the data, contacted CoD about terms of use
    • sci-arc meeting. Clarification for SHARK-VIS: once in Tucson, just a buffer for netwrok transfer to Italy.
    • ARGOS meeting, UAO/SOUL work needed
    • ROWG#6 settle on 42"+ configs, 2 flats, 1 curved 49" less pixels - draft remote operation requirement document
    • Observatory planning restat, 3 options
  • This week
    • remote operation requirement document
    • DMS
      • weather loss, websockets, AdSec, etc.
      • LBC stiched png preview
    • SHARK-VIS new IIF document and clarify DD entries for SOUL
    • follow up student ad
    • cobbler / kickstart, ims f30

XZ

  • Last week
    • Finalize the PB action items template with Guido
    • Prepare merge UAO to SOUL, thanks to Yang and Matthieu for giving git information.
    • PB SW/HW
    • Meetings
  • This week
    • Start working on PB action items
    • Merge UAO to SOUL
    • Review the slopeBCU document
    • PB SW/HW

DE

  • Veeam license quote request (agent model for oVirt)
  • Looked at Autodesk licensing
    • From my perspective it looks fine
    • Hearing rumblings about licensure failings; investigating
  • Re-ran SMART test on Synology disk that was complaining; definitely showing pre-fail now. Still not critical, but we should replace it next time IT is in Steward
  • Student account creation for JH optics student
  • Tracked Robert down to investigate Vault client issue
    • Software got into a bad install state; genuinely did not work to the point that I couldn't uninstall it
    • Ketiv offered a possible solution in a MS force uninstaller tool
    • Watiting on Robert to have another remote session
  • Safety training
  • Veeam backup tweaks to skip deprecated VMs
  • Vibration laptop work and handoff to Peter
  • KVM troubleshooting ongoing
  • Send training request to orders@
  • LeRoy came through with good images of luci.luci
    • Confirmed that Luci interface cards are PCIe x1, good for future upgrades
    • Plenty of available slots for RAID card
    • Individual SATA cables into backplane, will use SFF-8087 --> 4x SATA adapter cable from RAID card
  • Campus network outage tomorrow (5 May) morning 0430-0730
    • "may" experience interruptions to campus resources
    • Given that we are WFH, it is unlikely to have significant impact on our users

YZ

  • Previous:
    • AO-Cravat Python code update and some investigation.
    • modsAlign (Python3 updating and xpaget refactoring).
    • MODS WCS header discussion with Olga (and Rick).
    • Follow up on TMS Mathematica questions (Got some new info from Andrew on weekend).
    • More safety training courses appeared!
  • This week:
    • More work on AO-Cravat (Python code update and start to look into document generation).
    • Continue modsAlign refactoring work.
    • Review Andrew's new TMS information and come up with some plan on TMS software.
    • Discuss with Olga on MODS WCS header plan.
    • Follow up on TCS/IIF VCAN document with Robert (he says he's having some trouble with VCAN).
    • AO daily training session.
    • Finish some safety training.

CH

Last Week:
    • DMS: updated description file generator, began related changes to imonnit ingest script.
This Week:
    • DMS: complete imonnit ingest changes.
    • OA Control: Research software changes needed for encoder support binary output instead of Gray code.

IS

Last week:
  • PIT
    • Worked on PIT admin tool (authorization)
  • OT
    • MODS twilights issues on scrip generation.
    • Fix on LUCI lamps xml. Luci requires to have a complete and populated lamps/lamp array. The fix was not to remove the lamps element but the extra observation item we use to insert in LUCI lamp calibrations
  • LUCI
    • Look more in detail on the GEIRS software.
    • Trying to get the LUCI server environment working in mac, thought of using docker but exposing RMI ports via docker might be an issue, will do a VM instead.

This week:
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • work with Andrew on current test OT release issues
  • LUCI
    • Get the LUCI server environment working in mac.
    • Once I get the previous point to work, get on #IT7952 an RTD/Aladin GUI issue.
    • Look into LUCI WebIO device configuration.

SH

  • re-wiring in Tucson computer room
    • moved abell, but cannot run it permanently
  • consolidate cveillet-shell on to shell
  • monitor/keyboard/mouse for Jim
  • network UPS Tucson
  • Cindy iPad

next:

  • campus network upgrade tomorrow
  • DNS and abell to ovirt
  • web camera access for MGIO
  • more Tucson wiring now that UITS is done
  • creation of new private network.
Topic revision: r14 - 04 May 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback